# thanks: Corresponding author

### v1

@@ Introduction
The structure of this paper is as follows. In Sec. I we formulate GEM coupling on Kerr spacetime within the Newman–Penrose formalism and derive explicit expressions for the quadratic electromagnetic source of the gravitational Teukolsky equation. In Sec. II we briefly recap the associated GEM QQNMs and their relation to linear gravitational and electromagnetic QNMs. In Sec. III we assess the plausibility of observing GEM effects in several astrophysically motivated scenarios. We summarize our findings and outline future directions in Sec. IV.
@@ I GEM in Kerr spacetime
In either case, the particular solution is obtained through convoluting the Green’s function—for the Teukolsky kernel [18]—of the electromagnetic perturbation with the effective source. The resulting solution is then used to construct the effective source for the gravitational perturbation via the electromagnetic stress-energy tensor [38, 37], allowing for the analysis of GEM modes at the first level of mixing. This is a concise sketch of the procedure. Here we only focus on the perturbed, minimally coupled Einstein–Maxwell system on the Kerr background and examine the first level of mixing [^1] The first level of mixing refers to the hierarchy between the energy scales of “pure” gravitational phenomena and electromagnetic ones. See [38, 37] for further discussion.
@@
As shown in [38], the $T_{4}^{\mathrm{EM}}$ can be rearranged to be decomposed into two pieces: $T^{\text{lin}}_{4}\left(\phi^{(1)}_{i},J^{\text{Matter}}\right)$ and $T^{\text{quad}}_{4}\left(\phi^{(1)}_{i}\right)$. The first term is primarily relevant in EMRI problems and controls GEM contributions that are linear in the electromagnetic perturbations, while the second term is quadratic. As discussed in this work and further analyzed in [38, 37], minimal coupling dictates that GEM effects naturally manifest as linear and quadratic in the electromagnetic perturbations.
@@
where $\Delta\equiv r^{2}-2Mr+a^{2},\qquad R^{2}\equiv r^{2}+a^{2}$. We can express the contribution of electromagnetic source terms to gravitational perturbations as follows:
@@ I.1 Constructing $\Phi_{1}$
$\begin{gathered}{}_{s}T\,{}_{s}R=0\quad\quad{}_{s}O\,{}_{s}S=0,\end{gathered}$ (10)
@@
$\begin{gathered}K=\omega R^{2}+am,\quad Q=a\omega\sin\theta+m\csc\theta,\end{gathered}$ (12)
@@ I.2 Gravity-Source Separation
The particular solutions to the cases mentioned earlier are expected to take a separable form. A practical approach involves leveraging the orthonormality of spin-weighted spheroidal harmonics [41].
@@ II GEM Coupling
After coalescence of BBH, the remnant black hole is believed to relax into a state described by the Kerr solution by emitting GWs characterized by quasinormal modes (QNMs) [43, 44, 45, 46, 47]. Accordingly, at higher-order calculations, the coupling between gravitational and electromagnetic QNMs is expected to induce a new quadratic QNM “flavor”, which we refer to as GEM QQNMs, emerging in both the gravitational and electromagnetic sectors [^2] A more rigorous definition of GEM QQNMs, as poles in frequency space, is provided in complementary work [38]..
Focusing on the outgoing radiative degrees of freedom—namely, the Weyl scalar $\psi^{\textit{GEM}}_{4}$ and the Maxwell scalar $\phi^{\textit{GEM}}_{2}$—the governing equations as we discussed above could be written as follows:
$_{-2}\mathcal{T}\psi_{4}^{\textit{(GEM)}}=T^{\textit{EM-EM}}\left(\phi^{(1)}_{i}\right)$, (15)
@@
where the source term $T^{\textit{G-EM}}$ includes only the cross terms of first order gravitational and electromagnetic perturbations, which requires reconstructing the metric [14, 15, 48, 49] along with the Maxwell tensor. In (15) and (16), the operator $_{s}\mathcal{T}$ denotes the Teukolsky operator for spin weight $s$.
In [37, 38], authors studied the GEM QQNMs frequencies for the both gravitational and electromagnetic sectors, respectively, that were given by
@@
$\Omega^{\textit{(GEM)}}_{(l_{1},m_{1},n_{1})\times(l_{2},m_{2},n_{2})}$ $=\Omega_{(l_{1},m_{1},n_{1})}\pm\omega_{(l_{2},m_{2},n_{2})}.$ (18)
Here, $\omega_{(l,m,n)}$ and $\Omega_{(l,m,n)}$ represent the gravitational and electromagnetic QNM frequencies, respectively.
@@
GEM QQNMs are expected to have longer lifetimes compared to the first overtone in each sector, as exemplified for the $(l=4,m=4)$ and $(l=2,m=2)$ harmonic in fig. 1. This behavior is qualitatively similar to that of GG QQNMs, reflecting the similarity of electromagnetic and gravitational QNMs. On the other hand, it is worth noting that GEM QQNMs are not solely dictated by their parent modes, as is the case for GG QQNMs, but also by the nature of the coupling between electromagnetic and gravitational fields, which is the minimal coupling in (17) and (18).
[…]
For instance, if instead of the minimally coupled Maxwell-Einstein system, we studied the Maxwell-Dilaton-Einstein system[47], the QNMs in (17) and (18) would differ, see fig 2. Such studies could provide valuable tools for testing the principle of minimal coupling and probing alternative theories of gravity, including scalar-tensor or dilaton-like extensions of general relativity.
Figure 2: This figure illustrates how the frequencies of the polar gravitational QNM $(2,0)$ and the polar GEM QQNM $(1,0)\times(1,0)$ are changing on a static charged background with a charge-to-mass ratio $Q/M=0.6$ in the context of the Maxwell-Dilaton-Einstein theory as parameterized by the dilaton coupling $\eta\in[0,1.52]$. The stars denote the corresponding QNMs predicted by the minimally coupled Maxwell-Einstein theory at $\eta=0$ for both modes. This figure serves as a proof of concept, demonstrating that modifying the coupling mechanism directly influences the QNM frequencies. As a result, it presents an intriguing opportunity to test the minimal coupling principle. The data presented in this figure are drawn from [47, 50].
@@ III Discussion
From this point forward, we focus on the gravitational sector and further examine (15) and (17). In the context of the ringdown analysis, the situation differs slightly from EMRIs, as defining a perturbation parameter is not straightforward in general [51]. On the other hand, numerical simulations of head-on or quasi-circular inspirals of charged BBHs [52, 53, 54, 55, 56, 57, 58] highlight that both gravitational and electromagnetic radiation are emitted during all phases of coalescence. The ringdown phase, as expected, is characterized by QNM frequencies from both types of radiation. Depending on the binary’s parameters, electromagnetic contributions can represent a significant fraction of the total emission, underscoring their importance in black hole spectroscopy for this category of mergers[43, 59, 60].
Charged black hole mergers can involve two like-charged holes, one charged and one neutral, or oppositely charged components. The charge configuration critically shapes the dynamics and radiation of the remnant. In the oppositely charged case, conservation laws imply a net charge $Q_{\text{tot}}=||Q_{1}|-|Q_{2}||$, and energy is radiated primarily through low multipole modes, notably the dipole $(\ell=1,m=1)$. In such scenarios, electromagnetic (EM) radiation can rival or exceed gravitational radiation, making nonlinear GEM interactions as significant as purely gravitational (GG) ones. Indeed, for head-on collisions, it was shown in [53] that EM dipole emission becomes comparable to gravitational waves for $|Q|/M\gtrsim 0.2$, and even dominates for $|Q|/M\gtrsim 0.37$. Similar behavior is expected in more general merger configurations.
To investigate nonlinear effects in numerical simulations, one typically uses envelope-based estimates to assess the plausibility of identifying quadratic signatures. For example, one can compare the noise level in waveforms—extracted at null or spatial infinity—to the square of the parent mode waveform, in order to judge whether QQNMs could manifest in child modes. This approach applies to both GG and GEM interactions. Moreover, if gravitational radiation is expected to dominate certain harmonics, we can directly assess whether GEM contributions are comparable to those of GG interaction by comparing the energy radiated in both electromagnetic and gravitational channels. This offers a practical criterion for determining whether GEM QQNMs can be meaningfully extracted against the noise estimated in the child gravity mode. Numerous studies [7, 8, 11, 61, 62, 63, 64] have explored the correlation between the amplitudes of GG QQNMs, $\psi_{4}^{(2)}$, and their parent mode amplitudes, $\psi_{4}^{(1)}$, as well as how this relationship can depend on spin, initial data, and mode parity. This correlation is typically expressed through the ratio:
@@
Suppressing the lmnlmn indices,
$\Gamma\sim\frac{B^{(1)2}}{A^{(1)2}}=|\omega|\frac{\dot{E}^{RD}_{EM}}{\dot{E}^{RD}_{GW}},$ (22)
where $|\omega|=\omega\omega^{*}$, and $\dot{E}^{RD}$ is the rate of energy emission during the ringdown phase. For instance, in the opposite charge binaries, $\Gamma$ magnitude should be greater than in same charge systems.
@@
Qualitatively, GEM effects can emerge during the inspiral phase of an EMRI due to the energy content of the electromagnetic field [^3] The situation can become even more intriguing if QQNMs can be identified in the inspiral signal, akin to the case with linear QNMs in [73]. . Sophisticated calculations within the framework of post-Newtonian formalism or black hole perturbation theory could provide hints whether GEM effects could be detected in signals relevant to the upcoming LISA mission. Although we aim to address such calculations in future work, here we take the first step by assessing the conditions under which GEM effects become significant in second-order calculations using a simple energy comparison argument.
We consider a generic scenario involving a neutron star (NS) or magnetar of mass $m$ inspiraling around a supermassive black hole (SMBH) of mass $M$. Our goal is to determine when the electromagnetic energy contributes non-negligibly to waveform modeling compared to purely gravitational effects. Taking advantage of the nature of the system, we use the small quantities $m/M$ and $U_{EM}/Mc^{2}$ as proper perturbation parameters. Consider the NS with a mass ranging from $1-2\,M_{\odot}$ and a radius of about $10$ km, surrounded by a magnetic field with a strength between $10^{7}$ and $10^{12}\,\mathrm{T}$. If this magnetic field extends outward to a region approximately 121-2 times the neutron star’s radius, a rough estimate of the electromagnetic energy content in the surrounding region will be
$\frac{U_{EM}}{Mc^{2}}\sim\frac{10^{-13}M_{\odot}}{M}\,\text{to}\,\frac{10^{-3}M_{\odot}}{M},$ (23)
where $U_{EM}=B^{2}/2\mu_{0}$, while for the mass ratio we used
$\frac{m}{M}\sim\frac{2M_{\odot}}{M}.$ (24)
For systems with total mass $M_{\text{tot}}\geq 10^{6}\,M_{\odot}$, the terms $(m/M)^{2}$ and $U_{EM}/Mc^{2}$ become comparable at the lower limit of magnetic field strength $B\sim 10^{7}\,\mathrm{T}$ (which is the most pessimistic case). Naturally, more strongly magnetized NSs would amplify the GEM effects, making them relevant even in systems with lower $M_{\text{tot}}$. Hence, neglecting the GEM effects in the second-order calculations for NS-SMBH systems could introduce biases in gravitational waveform analysis. As EMRIs are one of the main sources for LISA, this estimate may warrant further investigation in this direction.
@@ IV Conclusion
In this work we have extended previous studies of GEM coupling on Schwarzschild backgrounds to the more astrophysically relevant case of Kerr spacetime. Within Teukolsky formalisms, we derived explicit quadratic electromagnetic sources for the gravitational Teukolsky equation and re-visited how GEM interactions generate a new family of GEM QQNMs in both the gravitational and electromagnetic sectors. These modes enrich the spectrum beyond purely gravitational quadratic QNMs and provide a concrete framework in which to organize nonlinear GEM effects in rotating black hole spacetimes.
We then examined several astrophysical situations in which GEM signatures may become comparable to, or even rival, purely gravitational nonlinearities. For charged BBH mergers, existing simulations indicate that electromagnetic radiation can contribute a non-negligible fraction of the radiated energy, especially in oppositely charged configurations, suggesting that GEM QQNMs could be detectable under the same conditions where quadratic gravitational modes are already accessible. In magnetized mergers such as BNS or NSBH systems, the presence of ultra-strong magnetic fields, potentially as large as $B\sim 10^{12}\,\mathrm{T}$, raises the prospect that GEM effects may leave measurable imprints in the ringdown of the remnant, although the non-vacuum nature of these environments complicates the analysis. For EMRIs, simple energetic estimates indicate that the electromagnetic energy content of a strongly magnetized secondary can compete with purely gravitational second-order effects in parts of the parameter space relevant for LISA, implying that neglecting GEM contributions could bias waveform modeling.
Beyond waveform systematics, GEM modes open a number of conceptual and phenomenological opportunities. Because any form of energy gravitates, GEM signatures provide a natural handle on hidden sectors, including black holes charged under a dark $U(1)$ gauge symmetry that is otherwise decoupled from the Standard Model. Moreover, modifying the coupling between gravity and electromagnetism—for instance in Maxwell–dilaton–Einstein or scalar–vector–tensor theories—alters the GEM QQNM spectrum and offers a concrete way to test the minimal coupling principle. While magnetized mergers are more common in nature, charged BBH mergers constitute a particularly clean theoretical laboratory in which to benchmark GEM effects. A natural next step is therefore to implement the Kerr GEM sources derived here in perturbative and numerical frameworks, apply them to existing and future simulations of charged and magnetized systems, and map out the regions of parameter space where GEM signatures become observable with ground-based detectors and LISA.
@@
We extend our deepest gratitude to Mahmoud A. Mansour for his invaluable time, support, and assistance throughout this project. We also sincerely appreciate the insightful conversations with Jacob Fields and Jaime Redondo. Additionally, we are grateful to Vasileios Paschalidis, David Radice, and Elias Most for their valuable discussions. Finally, we acknowledge the partial support of D.S. through the U.S. National Science Foundation under Grant No. PHY-2310363.

### v2

@@ Introduction
The structure of this paper is as follows. In Sec. I we formulate GEM coupling on Kerr spacetime within the Newman–Penrose formalism and derive explicit expressions for the quadratic electromagnetic source of the gravitational Teukolsky equation. In Sec. LABEL:GEM_Coupling we briefly recap the associated GEM QQNMs and their relation to linear gravitational and electromagnetic QNMs. In Sec. III we assess the plausibility of observing GEM effects in several astrophysically motivated scenarios. We summarize our findings and outline future directions in Sec. IV.
@@ I GEM in Kerr spacetime
In either case, the particular solution is obtained through convoluting the Green’s function—for the Teukolsky kernel [18]—of the electromagnetic perturbation with the effective source. The resulting solution is then used to construct the effective source for the gravitational perturbation via the electromagnetic stress-energy tensor [38, 37], allowing for the analysis of GEM modes at the first level of mixing [^1] The first level of mixing refers to the hierarchy between the energy scales of “pure” gravitational phenomena and electromagnetic ones. See [38, 37] for further discussion.. This is a concise sketch of the procedure. Here we only focus on the perturbed, minimally coupled Einstein–Maxwell system on the Kerr background.
@@
As shown in [38], the $T_{4}^{\mathrm{EM}}$ can be rearranged to be decomposed into two pieces: $T^{\text{lin}}_{4}\left(\phi^{(1)}_{i},J^{\text{Matter}}\right)$ and $T^{\text{quad}}_{4}\left(\phi^{(1)}_{i}\phi^{(1)}_{j}\right)$. The first term is primarily relevant in EMRI problems and controls GEM contributions that are linear in the electromagnetic perturbations, while the second term is quadratic. As discussed in this work and further analyzed in [38, 37], minimal coupling partially control how GEM effects manifest as linear and quadratic in the electromagnetic perturbations.
@@
where $|\Phi_{2}^{(1)}|^{2}=\Phi_{2}^{(1)}\Phi_{2}^{(1)\,*}$, $\Delta\equiv r^{2}-2Mr+a^{2},\qquad R^{2}\equiv r^{2}+a^{2}$. We can express the contribution of electromagnetic source terms to gravitational perturbations as follows:
@@ I.1 Constructing $\Phi_{1}$
[…]
where $_{s}R$ and $_{s}S$ are the Teukolsky radial and angular functions and $s=\pm 1$. Subsequently, perturbed Maxwell scalars could be written as
@@
$\begin{gathered}K=\omega R^{2}+am,\\ Q=a\omega\sin\theta+m\csc\theta,\end{gathered}$ (12)
@@ I.2 Gravity-Source Separation
The particular solution to the cases mentioned earlier are expected to take a separable form. A practical approach involves leveraging the orthonormality of spin-weighted spheroidal harmonics [41].
@@ II GEM Coupling
After the coalescence of a BBH, the remnant black hole is believed to relax into a state described by the Kerr solution by emitting GWs characterized by QNMs [43, 44, 45, 46, 47]. Accordingly, at higher orders in perturbation theory, the coupling between gravitational and electromagnetic QNMs is expected to induce a GEM QQNM “flavor”. [^2] A more rigorous definition of GEM QQNMs, as poles in frequency space, is provided in complementary work [38].
Focusing on $\psi^{\textit{(GEM)}}_{4}$ and $\phi^{\textit{(GEM)}}_{2}$, the governing equations can be written as
$_{-2}\mathcal{T}\psi_{4}^{\textit{(GEM)}}=T^{\textit{EM-EM}}\!\left(\phi^{(1)}_{i}\phi^{(1)}_{j}\right)$, (15)
@@
where the source term $T^{\textit{G-EM}}$ includes only the cross-terms of first-order gravitational and electromagnetic perturbations, which requires reconstructing the metric [14, 15, 48, 49] along with the Maxwell tensor. In (15) and (16), the operator $_{s}\mathcal{T}$ denotes the Teukolsky operator for spin weight $s$.
In [37, 38], the GEM QQNM frequencies for the gravitational and electromagnetic sectors were found to be
@@
where $\omega_{(l,m,n)}$ and $\Omega_{(l,m,n)}$ represent the gravitational and electromagnetic QNM frequencies, respectively.
@@
GEM QQNMs are expected to have longer lifetimes compared to the first overtone in each sector, as exemplified for the $(l=4,m=0)$ and $(l=2,m=0)$ harmonics in Fig. 1. This behavior is qualitatively similar to that of GG QQNMs, reflecting the similarity of electromagnetic and gravitational QNMs.
Moreover, GEM QQNMs are not solely dictated by their parent modes, but also by the underlying coupling between the electromagnetic and gravitational sectors. In our minimally coupled Einstein–Maxwell setup this coupling is fixed by the standard stress–energy source terms entering Eqs. (17) and (18). In this sense, the GEM amplitudes in this framework probe essentially the same interaction structure that sources the linear electromagnetic and gravitational modes, so any departure from minimal coupling would manifest itself as a deformation of the associated spectra.
A useful example of this sensitivity is provided by Einstein–Maxwell–dilaton (EMD) theory [47]. Pacilio and Brito computed the gravitational and electromagnetic QNM spectra of weakly charged, static and slowly rotating EMD black holes as functions of the dilaton coupling $\eta$ and charge-to-mass ratio $v=Q/M$. They found that, in the static case, the fundamental gravitational QNMs depend only weakly on $\eta$ (with relative shifts at the $\lesssim\mathcal{O}(10^{-2})$ level across $0\lesssim\eta\lesssim 3$), whereas the electromagnetic spectrum exhibits a much stronger dilaton dependence and a pronounced breaking of axial/polar isospectrality, reaching $\sim 10\%$ differences in the polar sector for $v\sim 0.6$ [47].
In Fig. 2 we exploit their tabulated static QNMs for $v=Q/M=0.6$ to construct the corresponding polar GEM QQNM $(1,0)\times(1,0)$ from the quadratic combination of the parent polar electromagnetic mode $(1,0)$, and to compare it with the ordinary polar gravitational mode $(2,0)$, both shown as functions of the dilaton coupling $\eta$. While the gravitational mode frequency varies only mildly with $\eta$, the electromagnetic parent mode is significantly more $\eta$-sensitive, so that its quadratic GEM descendant inherits an even larger fractional variation than the linear gravitational spectrum. This proof-of-concept example suggests that, in EMD-like theories and, more generally, in scalar–tensor or tensor–vector–scalar extensions where the electromagnetic sector is coupled to additional fields, GEM QQNMs can be parametrically more sensitive to the coupling than both linear gravitational QNMs and GG quadratic modes at the level of the underlying spectrum.
This pattern is broadly consistent with other families of gravity–electromagnetism couplings studied in the literature. Nonminimal Einstein–Maxwell models with curvature couplings of the schematic form $RF^{2}$ or $R_{\mu\nu}F^{\mu\alpha}F^{\nu}{}_{\alpha}$ modify the propagation of electromagnetic fields on curved backgrounds and are expected to deform the electromagnetic-led QNM spectra more strongly than the tensor sector [50, 51]. Similarly, in Einstein–Maxwell–scalar theories with couplings $f(\phi)F^{2}$, scalarized charged black holes develop mixed scalar/electromagnetic/gravitational modes whose frequencies and damping times can be highly sensitive to the scalar–Maxwell coupling function [52, 53, 54], and in which axial–polar isospectrality is generically broken in the electromagnetic and gravitational sectors [53]. In all these cases the electromagnetic (or scalar–electromagnetic) channels typically respond more strongly to changes in the coupling than the purely tensorial sector.
Figure 2: Frequencies of the polar gravitational QNM $(2,0)$ and the polar GEM QQNM $(1,0)\times(1,0)$ on a static charged background with charge-to-mass ratio $v=Q/M=0.6$ in Einstein–Maxwell–dilaton theory, shown as functions of the dilaton coupling $\eta\in[0,1.52]$. The stars denote the corresponding QNMs predicted by minimally coupled Einstein–Maxwell theory at $\eta=0$ for both modes. The electromagnetic sector exhibits a much stronger $\eta$-dependence than the gravitational one, so that the GEM quadratic mode inherits an enhanced sensitivity to the dilaton coupling at the level of the spectrum. This figure is intended as a proof of concept for the coupling dependence of GEM QQNMs; whether such differences are observationally accessible depends on the achievable signal-to-noise ratio, detector sensitivity, and systematic uncertainties, and we do not attempt a detailed forecast here. The data used in this figure are taken from [47, 55].
@@ III Discussion
In the context of the ringdown analysis, the situation differs slightly from EMRIs, as defining a perturbation parameter is not straightforward in general [56]. Yet, numerical simulations of head-on or quasi-circular inspirals of charged BBHs [57, 58, 59, 60, 61, 62, 63] highlight that both gravitational and electromagnetic radiation are emitted during all phases of coalescence. The ringdown phase, as expected, is characterized by QNM frequencies from both types of radiation. Depending on the binary’s parameters, electromagnetic contributions can represent a significant fraction of the total emission, underscoring their importance in black hole spectroscopy for this category of mergers[43, 64, 65].
III.1 Charged black hole mergers
Charged black hole mergers can involve two like-charged holes, one charged and one neutral, or oppositely charged components. The charge configuration critically shapes the dynamics and radiation of the remnant. In the oppositely charged case, conservation laws imply a net charge $Q_{\text{tot}}=||Q_{1}|-|Q_{2}||$, and energy is radiated primarily through low multipole modes, notably the dipole $(\ell=1,m=1)$. In such scenarios, electromagnetic radiation can rival or exceed gravitational radiation, making nonlinear GEM interactions as relevant as GG. Indeed, for head-on collisions, it was shown in [58] that EM dipole emission becomes comparable to gravitational waves for $|Q|/M\gtrsim 0.2$, and even dominates for $|Q|/M\gtrsim 0.37$. Similar behavior is expected in more general merger configurations.
To investigate nonlinear effects in numerical simulations, one typically uses envelope-based estimates to assess the plausibility of identifying quadratic signatures. For example, one can compare the noise level in waveforms—extracted at null or spatial infinity—to the square of the parent mode waveform, in order to judge whether QQNMs could manifest. This approach applies to both GG and GEM interactions. Moreover, if gravitational radiation is expected to dominate certain harmonics, we can directly assess whether GEM contributions are comparable to those of GG interaction by comparing the energy radiated in both electromagnetic and gravitational channels. This offers a first-step, practical criterion for determining whether GEM can be meaningfully extracted against the noise estimated in the child gravity mode. Numerous studies [7, 8, 11, 66, 67, 68, 69] have explored the correlation between the amplitudes of GG QQNMs, $\psi_{4}^{(2)}$, and their parent mode amplitudes, $\psi_{4}^{(1)}$, as well as how this relationship can depend on spin, initial data, and mode parity. This correlation is typically expressed through the ratio:
@@ III.1 Charged black hole mergers
To leading order at infinity,
$\begin{gathered}\psi_{lmn\,4}^{(1)}\simeq A_{lmn}^{(1)}{}_{-2}S_{lm}\,e^{i\omega(t-r_{*})+im\varphi},\\ \phi_{lmn\,2}^{(1)}\simeq B_{lmn}^{(1)}{}_{-1}S_{lm}\,e^{i\omega(t-r_{*})+im\varphi},\end{gathered}$ (22)
with the spin–weighted spheroidal harmonics normalized. The gravitational and electromagnetic powers carried by this single mode are
$P_{\rm GW}\;=\;\frac{|B^{(1)}|^{2}}{4\pi\,|\omega|^{2}},\qquad P_{EM}\;=\;\frac{|A^{(1)}|^{2}}{4\pi}.$ (23)
as we focusing on dominate mode of each chanel, the by suppressing the lmnlmn indices,
$\Gamma\sim\frac{B^{(1)2}}{A^{(1)2}}=|\omega|^{2}\frac{\dot{E}^{RD}_{EM}}{\dot{E}^{RD}_{GW}},$ (24)
where $\dot{E}^{RD}$ is the rate of energy emission during the ringdown phase. For instance, in the opposite charge binaries, $\Gamma$ magnitude should be greater than in same charge systems.
@@
Qualitatively, GEM effects can emerge during the inspiral phase of an EMRI due to the energy content of the electromagnetic field. Sophisticated calculations within the framework of post-Newtonian formalism or black hole perturbation theory could provide hints whether GEM effects could be detected in signals relevant to the upcoming LISA mission. We aim to address such calculations in future work.
@@ IV Conclusion
In this work we have extended previous studies of GEM coupling on Schwarzschild backgrounds to the more astrophysically relevant case of Kerr spacetime. Within Teukolsky formalisms, we derived explicit quadratic electromagnetic sources for the gravitational Teukolsky equation and re-visited how GEM interactions generate a new family of GEM QQNMs in both the gravitational and electromagnetic sectors. These modes enrich the spectrum beyond purely gravitational quadratic QNMs and provide a concrete framework in which to organize nonlinear GEM effects in rotating black hole spacetimes. For the charged BBH mergers, we expect the remannt to be charged hence, next stop in the analytical research should be the Kerr-Newman spacetime.
We then examined several astrophysical situations in which GEM signatures may become comparable to, or even rival, purely gravitational nonlinearities. For charged BBH mergers, existing simulations indicate that electromagnetic radiation can contribute a non-negligible fraction of the radiated energy, especially in oppositely charged configurations, suggesting that GEM QQNMs could be detectable under the same conditions where quadratic gravitational modes are already accessible. In magnetized mergers such as BNS or NSBH systems, the presence of ultra-strong magnetic fields, potentially as large as $B\sim 10^{12}\,\mathrm{T}$, raises the prospect that GEM effects may leave measurable imprints in the ringdown of the remnant, although the non-vacuum nature of these environments complicates the analysis.
Beyond waveform systematics, GEM modes open a number of conceptual and phenomenological opportunities. Because any form of energy gravitates, GEM signatures provide a natural handle on hidden sectors, including black holes charged under a dark $U(1)$ gauge symmetry that is otherwise decoupled from the Standard Model. Moreover, modifying the coupling between gravity and electromagnetism—for instance in Maxwell–dilaton–Einstein or scalar–vector–tensor theories—alters the GEM QQNM spectrum and offers a concrete way to test the minimal coupling principle. While magnetized mergers are more common in nature, charged BBH mergers constitute a particularly clean theoretical laboratory in which to benchmark GEM effects.
Finally, although we have focused on Kerr backgrounds as a clean starting point, charged mergers naturally point toward Kerr–Newman geometry [70, 71, 72, 40, 73, 74]. A natural next step is therefore to generalize the GEM analysis to the later one, implement it in perturbative solvers and numerical-relativity frameworks, and apply it to existing and future simulations of charged and magnetized systems. Mapping out the regions of parameter space where GEM signatures become comparable to GG nonlinearities, and assessing under what conditions they might be constrained with ground-based detectors and LISA, are key goals for future work.
@@
We extend our deepest gratitude to Mahmoud A. Mansour for his invaluable time, support, and assistance throughout this project. We also sincerely appreciate the insightful conversations with Jacob Fields and Jaime Redondo Yuste. Additionally, we are grateful to Vasileios Paschalidis, David Radice, and Elias Most for their valuable discussions. Finally, we acknowledge the partial support of D.S. through the U.S. National Science Foundation under Grant No. PHY-2310363.